OpenStudy (osanseviero):

How to demostrate that z+(-z)=0 in complex numbers

OpenStudy (osanseviero):

I've done two things: a +bi+ (-a -bi)= 0 And z-z+z=0+z a+bi=a+bi+0 a+bi+a+bi

OpenStudy (osanseviero):

I think I need something else